OpenStudy (anonymous):

Simplify the expression . 36 ^ -1/2

OpenStudy (anonymous):

negative exponent = 1 / positive

OpenStudy (anonymous):

so 1/(36^1/2)

OpenStudy (anonymous):

right - now you can simplify some more

OpenStudy (anonymous):

whats the fractional exponent mean?

OpenStudy (anonymous):

4 ^ 1/2 = 2 - that shoudl help

OpenStudy (anonymous):

exponent 1/2 = square root, 1/3 = cube root etc answer is 1 / sqrt 36 = 1/6